OpenStudy (anonymous):

Solve the equation. 8x-12=4x+24

OpenStudy (whpalmer4):

Subtract 4x from both sides, add 12 to both sides, can you see what to do from there?

OpenStudy (jhannybean):

First you want to get your x's all on one side, so here you can subtract 4x on both sides 8x - 12 - 4x = 4x - 4x + 24

OpenStudy (jhannybean):

simplify the equation 4x - 12 = +24 now isolate the x's on one side of the equation, and your constants on the other by adding 12 to both sides 4x - 12 + 12 = 24 + 12 simplify again 4x = 36 now divide 4 on both sides 4x/4 = 36/4 x = ?

OpenStudy (whpalmer4):

Just to illustrate that there's often no one set way to solve an equation: \[8x-12=4x+24\]Divide both sides by 4 \[2x-3=x+6\]Subtract \(x\) from both sides \[x-3=6\]I think you probably don't need my help at this point :-)